OpenStudy (cenaida):

Solve the system, using substitution. What is the n coordinate of the solution? 2m - 3n = 6 and m = 3 - 2n A. 2 B. 0 C. -4 D. 2 E. no solution

OpenStudy (anonymous):

do u want me to show u how to set it up or how to solve it ?

OpenStudy (cenaida):

yes please

OpenStudy (anonymous):

ok ;) the impoot is 2m - 3n = 6,m = 3 - 2n so ur job is to solv for n u will hve to find a way to get n bye its self, when u do that it will end up being m=1-n every letter equals 1 in some equations what u want to do next is switch m and n then the equation will look like n=1-m, (m=1) so whats the answer ?